OpenStudy (anonymous):

Fast Please answer Algebra 2 How do I do. Find a quadratic equation with integral coefficients having the given roots. 1+ √3, 1 - √3

OpenStudy (anonymous):

I got a[x^2-2x-2=0 is that correct?

OpenStudy (agent0smith):

If a quadratic has roots a and b, you can write it as (x - a)(x - b) Your roots are (1+ √3) and (1 - √3) so you just have to enter those in place of a and b (and FOIL/expand if necessary) and that should satisfy the question.

OpenStudy (anonymous):

so can i skip the step of a[x^2- (sum of roots)x + (product of roots)]=0 step?

OpenStudy (anonymous):

or is it the same thing as (x-a)(x-b) but written diffrently?

OpenStudy (agent0smith):

You could probably do it that way too... should work.

OpenStudy (anonymous):

they are correct. or the back of the back says so.

OpenStudy (agent0smith):

Yeah, my mistake, they are.

OpenStudy (anonymous):

one quick question, does (x-b)(x-a) always work?

OpenStudy (agent0smith):

Yes, it should. That's where the roots come from. \[ax^2-2x-2 \]looks good to me.

OpenStudy (agent0smith):

Even if the roots are complex numbers, you can put them into (x-b)(x-a)

OpenStudy (anonymous):

As 2times 3 = 3times2, (x-a)(x-b) = (x-b)(x-a)